Conslder a largo population or interest its distributIon Is normal and ts mean 185 and standard deviatlon is 1 10. Let X-a alnglC 0030~atlon Round al probabilties to four decimals) a) Find PO190 b) Find FiX 190 Conelder anothr large populaton o interest ITs distntution is unknoun and its mean Is 185 and tstandand davaton is 110. Let-a snale observaton
Answer #1

a)

for normal distribution z score =(X-μ)/σ
here mean=       μ= 185
std deviation   =σ= 110.0000
probability = P(X<190) = P(Z<0.05)= 0.5199

b)

sample size       =n= 130
std error=σ=σ/√n= 9.6476
probability = P(Xbar<190) = P(Z<0.52)= 0.6985

c)

as sample size is greater then 30 therefore sample mean can be apprximated as normally distributed

hence

probability = P(Ybar<190) = P(Z<0.52)= 0.6985